v/8 = 2
the solution is 16. please help!
thank you!

Answer by oberobic(2304)

v/8 = 2
.
multiply both sides by 8 to eliminate the fraction
.
8*(v/8) = 8*2
.
the 8's cancel on the left hand side
.
v = 16
.
Done.